# A new perspective on the equivalence between Weak and Strong Spatial Mixing in two dimensions

### v1

@@ 1.1 Mixing in lattice models
This stronger notion is extremely powerful: indeed, it has a wealth of equivalent formulations in terms of analytic properties of the model, existence of expansions for correlation functions, and fast relaxation of the Glauber dynamic associated to the model, see [5, 6, 7, 13, 15]. This notion is relatively easy to establish in many models under a “very high temperature” assumption, using cluster expansion. Non-perturbative results are much harder to obtain, and usually require a lot of additional structure. Two famous cases are: 1) the monomer-dimer model on $\mathbb{Z}^{d}$[14], and 2) the Ising model on $\mathbb{Z}^{d}$ above the critical temperature, [3].
A very special case is the case of dimension two. Indeed, in this case, the boundary of a (reasonable) system is one-dimensional, and one dimensional systems famously have systematic exponential decay of every kind of correlations one can think of. So, it is natural to expect that systems satisfying weak mixing (exponential decay of information in the bulk) in dimension two (with a one-dimensional boundary left to perform information transfer) also satisfy the strong mixing condition. This picture was established in a large family of systems (finite range Gibbsian specifications) in [10], based on the dynamical criterion of [13]. Strong mixing for more particular models in dimension two under an “off-criticality assumption” was established in [12] (Ising models, 28 years before [3]), and [2] (Potts models, and FK-percolation). In the case of FK percolation, one additionally has to be careful with what is meant by “strong mixing”, and by “one-dimensional boundary conditions”, as FK-percolation has non-local weights (see the discussion in the Introduction of [2]).
@@
Sections 34, and 5 respectively give applications of the main result to finite range Gibbsian specifications, finite range FK percolation on $\mathbb{Z}^{2}$, and hard core models on $\mathbb{Z}^{2}$. Application to the Random Cluster representation of the Ashkin-Teller model is discussed in [4].
@@ 1.2 Notation, constants, surrounding sets
When not mentioned otherwise, $\mathbb{Z}^{2}$ and its subsets are endowed with the nearest-neighbour (with respect to the Euclidean distance) graph structure, connectivity and paths are with respect to that graph structure (edges are of the form $\{x,y\}$ with $|x-y|=1$). The connectivity and paths induced by the graph structure given by the nearest-neighbour graph with respect to $\mathrm{d}_{\infty}$ (edges are of the form $\{x,y\}$ with $\left\|x-y\right\|_{\scriptscriptstyle\infty}=1$) are called *-connectivity and *-paths. This notions are transported to $\mathbb{L}_{K}$ via the natural bijection between $\mathbb{Z}^{2}$ and $\mathbb{L}_{K}$ ($x\mapsto(2K+1)x$).
@@ 1.4 Hypotheses
If the model has the Markov property for the nearest-neighbour graph structure and a (weak form of) finite energy, Hypotheses (Mar1)(Mar2), and (Mar3) are immediately satisfied for any $l\geq 0$ (with $p_{\mathrm{bulk}}=1$). The same is true for range-one models (Markov property for the *-graph structure).
@@ 1.5 Main Result
Let $(\nu_{\Lambda}^{t})_{\Lambda,t}$ be as described in Section 1.3. Suppose that the mixing Hypotheses (Mix1), and (Mix2) with $C_{\mathrm{mix}}\geq 0,c_{\mathrm{mix}}>0$ hold. There is $p_{0}=p_{0}(c_{\mathrm{mix}})\in(0,1)$ such that if, for some $l\geq 0$, Hypotheses (Mar1)(Mar2) with $p_{\mathrm{bulk}}\geq p_{0}$, and (Mar3) with $\theta>0$ all hold, then, for any $p>0$, there is $\ell\geq 1$, and $q\in(0,1)$ such that for any $\Lambda\Subset\mathbb{Z}^{2}$, $t\in T$, $\Delta_{1},\Delta_{2}\subset\Lambda$, and $\xi,\xi^{\prime}\in\Omega_{\Delta_{1}}$ with $\nu_{\Lambda}^{t}(\sigma_{\Delta_{2}}=\xi)\nu_{\Lambda}^{t}(\sigma_{\Delta_{2}}=\xi^{\prime})>0$,
@@
$p_{i}=p_{i}(\Lambda,\Delta_{2})=\begin{cases}1&\text{ if }\Lambda_{\ell}(i)\cap\Delta_{2}\neq\varnothing,\\ p&\text{ if }\Lambda_{\ell+\ell^{\prime}}(i)\subset\Lambda\setminus\Delta_{2},\\ 0&\text{ if }\Lambda_{\ell}(i)\subset\Lambda^{c},\\ q&\text{ else }.\end{cases}$
This result is obtained from a slightly finer upper bound, but I do not believe that the finer result has a real added value compare to this one, and this statement is easier to digest. I also believe that a higher dimensional version of this result should hold (indeed: there the lack of strong mixing is manifested by the percolation of information along the boundary, which has arbitrarily large percolation parameter).
Strong mixing then follows for nice enough volumes: the percolation model $P_{\Lambda;\ell,\bar{p}}$ is very subcritical in the bulk and, when the system has a “nicely one-dimensional boundary”, the inhomogeneities close to the boundary cannot destroy this decay. In the applications, the case of large squares is considered as it is easy to show exponential decay there. But the result also holds with more complicated volumes. As far as applying the result is concerned, I believe Theorem 1.1 to be more convenient than putting some a priori restriction on the class of volume considered, and stating exponential decay for that class. Volumes under consideration might vary a great deal depending on what one is trying to do. Compared to the class of volumes considered in [2], which give strong mixing for simply connected volumes in the case of planar FK percolation, the present criterion is a priori neither weaker nor stronger: indeed, it is easy to find not simply connected regions for with it is easy to establish exponential decay under $P_{\Lambda;\ell,\bar{p}}$, and therefore apply Theorem 1.1, while it is not clear (and maybe wrong!) that $P_{\Lambda;\ell,\bar{p}}$ has exponential decay of connectivities uniformly over simply connected $\Lambda$. For the reader interested in non-trivial applications of Theorem 1.1, the arguments of [2, Section 2] permit to study exponential decay of connections probability under $P_{\Lambda;\ell,\bar{p}}$ for various family of volumes.
@@ 1.6 Relaxing the mixing hypotheses
Suppose moreover that Hypotheses (Mar1), and (Mar2) with $p_{\mathrm{bulk}}$ close enough to $1$ hold (independently of the value of $C_{*}\geq 0,c_{*}>0$). Then, Hypotheses (Mix1) and (Mix2) hold.
@@ 2.2 Patch exploration
Let $(\Omega,\mathcal{F},P)$ be a probability space. Let $I$ be a finite set, and $(\mathcal{X}_{x})_{x\in I}$ a collection of finite sets equipped with the discrete $\sigma$-algebra: $\mathcal{F}_{x}=\mathcal{P}(\mathcal{X}_{x})$. Set $\mathcal{X}=\bigtimes_{x\in I}\mathcal{X}_{x}$, $\mathcal{F}=\bigotimes_{x\in I}\mathcal{F}_{x}$, and let $X$ be a $\mathcal{X}$-valued random variable. Denote $X_{J}$ the restriction (projection) of $X$ to ×$\bigtimes_{x\in J}\mathcal{X}_{x}$. Suppose that for every $J\subset I$, and for $P$-almost every realization of $X_{J}$, $\xi$, the conditional measure $P(\cdot\,|\,X_{J}=\xi)$ is well defined. A patch exploration sampling for $X$ is a the data of
@@ 2.3 Coarse-Graining and rescaling
Let $l\geq 0$ be the parameter linked with the “Markov-type hypotheses”. It will be more convenient to work with the blocked field $X=\psi(\sigma)$. For the rest of the section, rescale everything by $12l+$ 1, and identify $\mathbb{L}_{l}$ with $\mathbb{Z}^{2}$, and $\mathbb{Z}^{2}$ with $(\delta_{l}\mathbb{Z})^{2}$, where $\delta_{l}=\tfrac{1}{2l+1}$. In particular, $\sigma$ lives on $(\delta_{l}\mathbb{Z})^{2}$ and $X$ lives on $\mathbb{Z}^{2}$. for $\Lambda\subset(\delta_{l}\mathbb{Z})^{2}$, introduce
@@
Note that they are related via: $\Theta_{ij}$ is the composition of $\Theta_{ji}$ with the mirror symmetry sending $\mathrm{e}_{ij}$ to $-\mathrm{e}_{ij}=\mathrm{e}_{ji}$. With these notations, and for $k\in\{\pm 1,\dots,\pm r^{L}\}$ let (see Figure 4 and Figure 5)
@@
Figure 6: Coarse-graining of a finite volume $\Lambda$: the small boxes are “blocked spins” ($\sigma$ lives on the disks, $X=\psi(\sigma)$ lives on the small boxes), the orange large boxes are $[\Lambda]_{\mathrm{bulk}}$, and the blue ones are $[\Lambda]_{\mathrm{bnd}}$.
@@ 2.4 Good configurations and canonical paths
Suppose that (Mar1), and (Mar2) with $p_{\mathrm{bulk}}$ close enough to $1$ hold. Then, there is $L_{0}\geq 1$, such that for any $L\geq L_{0}$, and any $i\in\Gamma$,
@@
[…]
@@
[…]
@@
Suppose that (Mar1), and (Mar2) with $p_{\mathrm{bulk}}$ close enough to $1$ hold. Then, for any $L\geq 1$, and any $e\in E_{\Gamma}$, $k\in\{\pm 1,\dots\pm r^{L}\}$,
@@
Proceed as in the proof of Claim 2: compare $\tilde{\mathcal{C}}_{E_{e}^{k}}$ with a Bernoulli percolation cluster (of parameter $p$ close to one when $p_{\mathrm{bulk}}$ is close to $1$). The absence of long crossing is then equivalent to the presence of short *-crossing of closed sites, which has exponential decay for $p$ large enough ($p$ larger than $161-e^{-2}/$ 16 is amply sufficient). ∎
@@ 2.6 Definition of the patch exploration process
Consider the planar labelled (multi)graph $G_{\ell}=(\Gamma,E_{\ell})$ ($\ell$ for labelled) with vertex set $\Gamma$ and $2r^{L}$ edges between neighbouring sites, labelled ±$\pm 1,\dots\pm r^{L}$, from the “inside” to the “outside” (see Figure 7)
@@
Now, the sequence of set valued functions $\mathcal{A}_{k},k\geq 0$ will be given by the simultaneous exploration of a percolation configuration and of a spin configuration. It is given by Algorithm 1, which requires a few preparations. Consider site-edges percolation configurations on $G_{\ell}$ with all sites of $[\Lambda]_{\mathrm{ext}}$ open, all edges with at least one endpoint in $[\Lambda]_{\mathrm{ext}}$ open, and all edges with both endpoints in $\Gamma\setminus[\Lambda]_{\mathrm{bulk}}$ open:
@@
  • no site in $[\Delta_{2}]$ is explored, and no edge with an endpoint in $[\Delta_{2}]$ is ever explored.
@@
  • $\mathcal{B}$ is the union of a collection of closed self-avoiding site-edge paths, $\gamma$, which surrounds $[\Delta_{2}]$ with some edges and sites in the set $[\Delta]_{2}$-surrounded by $\gamma$. $\eta_{\mathcal{B}}$ is open on $\gamma$;
@@ 2.7 Concluding the proof
$\sup_{\xi,\xi^{\prime}}\mathrm{d}_{\mathrm{TV}}\big{(}\nu_{\Lambda,\Lambda\setminus\Delta_{2}}^{t,\xi}|_{\Delta_{1}},\nu_{\Lambda,\Lambda\setminus\Delta_{2}}^{t,\xi^{\prime}}|_{\Delta_{1}}\big{)}\leq\mathrm{P}([\Delta_{1}]\text{ is not separated from }[\Delta_{2}]\text{ in }\tilde{\omega})$ (9)
@@
$q_{e}=\begin{cases}1-(1-\theta^{4L}/2)^{2r^{L}}\geq 1-e^{-2^{L}}&\text{ if }e\subset[\Lambda]_{\mathrm{bulk}}\\ \theta^{(2a+2)L^{2}}&\text{ if }e=\{i,j\},i\in[\Lambda],j\notin[\Lambda]_{\mathrm{bulk}}\\ 1&\text{ else }\end{cases}.$
@@ 3 Application I: Finite range Gibbsian specifications
Let $\Omega_{0}$ be a finite set. For $A\Subset\mathbb{Z}^{2}$, let $\phi_{A}:\Omega_{0}^{A}\to\mathbb{R}$ be a collection of uniformly bounded potential functions. The Gibbs specification associated to the collection $(\phi_{A})_{A\Subset\mathbb{Z}^{2}}$ is the collection of probability kernels $(\mu_{\Lambda}^{\xi})_{\Lambda\Subset\mathbb{Z}^{2},\xi\in\Omega_{0}^{\mathbb{Z}^{2}}}$ given by, for any event $A$ with compact support,
@@
where $Z_{\Lambda}^{\xi}$ is the normalization constant making $\mu_{\Lambda}^{\xi}$ a probability measure. Such a specification is called finite range when there is $R\geq 1$ such that $\phi_{A}=0$ for any $A$ with diameter larger than $R$.
@@
The main result of this section is a new proof of [10, Theorem 1.1].
@@
The set $(\mu_{\Lambda}^{\xi})_{\Lambda\Subset\mathbb{Z}^{2},\xi\in\Omega_{0}^{\mathbb{Z}^{2}}}$ is a particular case of the framework of Section 1.3. Take then $l>R$, the range of the potential. Then, the model satisfies (Mar1) with $\mathrm{Ma}_{u}=\Omega_{0}^{\Lambda_{3l+1}(u)}$, $\mathrm{Ma}_{u,\Lambda}^{\xi}=\Omega_{0}^{\Lambda_{3l+1}\cap\Lambda}$(Mar2) with $p_{\mathrm{bulk}}=1$, and (Mar3) with $\mathrm{a}_{u},\mathrm{a}_{u,\Lambda}^{\xi}$ any elements of the sets $\mathrm{Ma}_{u},\mathrm{Ma}_{u,\Lambda}^{\xi}$.
Now, $(\mu_{\Lambda}^{\xi})_{\Lambda,\xi}$ are assumed to be weak mixing. So, by Theorem 1.2, they satisfy Hypotheses (Mix1), and (Mix2). Let then $\ell,q$ be the length and percolation parameter guaranteed by Theorem 1.1 for $p=10^{-10}$. Then, for $L$ large enough compared to $\ell$ (depending on $q$), it is an exercise to show that the percolation models $P_{\Lambda;\ell,\bar{p}}$ have exponential decay of connectivity uniformly over $\Lambda$ a large enough square (see remark 3.1). ∎
@@
The statement is in fact more general: Theorem 1.1 compares mixing with connections in a (very) subcritical Bernoulli percolation with inhomogeneities. A sufficient condition for these model to have exponential decay should be that the inhomogeneity is "essentially one-dimensional" below a certain scale (e.g.: it is (well approximated by) a union of straight lines with large distances between intersection points, and lower bounded angles between lines). Obtaining optimal conditions under which such inhomogeneous site/block percolation has exponential decay seems like an interesting open problem. Handling “macroscopically one dimensional” inhomogeneities can be done as in [11, Lemma 3.2] (in particular, a similar argument applies in the large squares case of Theorem 3.1).

### v2

@@ 1.1 Mixing in lattice models
This stronger notion, also called complete analyticity (CA), is extremely powerful: indeed, it has a wealth of equivalent formulations in terms of analytic properties of the model, existence of expansions for correlation functions, and fast relaxation of the Glauber dynamic associated to the model, see [5, 6, 7, 15, 18]. This notion is relatively easy to establish in many models under a “very high temperature” assumption, using cluster expansion. Non-perturbative results are much harder to obtain, and usually require a lot of additional structure. Two famous cases are: 1) the monomer-dimer model on $\mathbb{Z}^{d}$, [16], and 2) the Ising model on $\mathbb{Z}^{d}$ above the critical temperature, [3].
A very special case is the case of dimension two. Indeed, in this case, the boundary of a (reasonable) system is one-dimensional, and one dimensional systems famously have systematic exponential decay of every kind of correlations one can think of. So, it is natural to expect that systems satisfying weak mixing (exponential decay of information in the bulk) in dimension two (with a one-dimensional boundary left to perform information transfer) also satisfy the strong mixing condition. Note that one has first to make sense of “reasonable system” to be able to expect anything: if a volume boundary is too fractal at the scale of the correlation length of the system (or even worse: at the scale of the interaction!), there is no hope that “weak implies strong” will hold generically. The relevant notion is called restricted complete analyticity (RCA): in terms of mixing, it asks that strong mixing holds for disjoints unions of sufficiently large squares. See [11, 12] for more details on this notion, and the discussion in the introduction of [17] for more on the failure of the full CA in models as nice as the Potts or Ising models on $\mathbb{Z}^{2}$ at values of the parameters for which weak mixing holds.
The equivalence between weak mixing and RCA was established in a large family of systems (finite range Gibbsian specifications) in [12], based on the dynamical criterion of [15]. Strong mixing for more particular models in dimension two under an “off-criticality assumption” was established in [14] (Ising models, 28 years before [3]), [17] (RCA for Potts model on $\mathbb{Z}^{2}$ with large enough $q$ above $T_{c}$), and [2] (Potts models, and FK-percolation on $\mathbb{Z}^{2}$, under a simply connected assumption on the volume). In the case of FK percolation, one additionally has to be careful with what is meant by “strong mixing”, and by “one-dimensional boundary conditions”, as FK-percolation has non-local weights (see the discussion in the Introduction of [2]).
@@
Sections 3, 4, and 5 respectively give applications of the main result to finite range Gibbsian specifications, finite range FK percolation on $\mathbb{Z}^{2}$, and hard core models on $\mathbb{Z}^{2}$. Application to the Random Cluster representation of the Ashkin-Teller model is discussed in [4].
@@ 1.2 Notation, constants, surrounding sets
When not mentioned otherwise, $\mathbb{Z}^{2}$ and its subsets are endowed with the nearest-neighbour (with respect to the Euclidean distance) graph structure, connectivity and paths are with respect to that graph structure (edges are of the form $\{x,y\}$ with |xy|=1|x-y|=1). The connectivity and paths induced by the graph structure given by the nearest-neighbour graph with respect to $\mathrm{d}_{\infty}$ (edges are of the form $\{x,y\}$ with $\left\|x-y\right\|_{\scriptscriptstyle\infty}=1$) are called *-connectivity and *-paths. This notions are transported to $\mathbb{L}_{K}$ via the natural bijection between $\mathbb{Z}^{2}$ and $\mathbb{L}_{K}$ ($x\mapsto(2K+1)x$).
@@ 1.4 Hypotheses
If the model has the Markov property for the nearest-neighbour graph structure and a (weak form of) finite energy, Hypotheses (Mar1), (Mar2), and (Mar3) are immediately satisfied for any $l\geq 0$ (with $p_{\mathrm{bulk}}=1$). The same is true for range-one models (Markov property for the *-graph structure).
@@ 1.5 Main Result
Let $(\nu_{\Lambda}^{t})_{\Lambda,t}$ be as described in Section 1.3. Suppose that the mixing Hypotheses (Mix1), and (Mix2) with $C_{\mathrm{mix}}\geq 0,c_{\mathrm{mix}}>0$ hold. There is $p_{0}=p_{0}(c_{\mathrm{mix}})\in(0,1)$ such that if, for some $l\geq 0$, Hypotheses (Mar1), (Mar2) with $p_{\mathrm{bulk}}\geq p_{0}$, and (Mar3) with $\theta>0$ all hold, then, for any $p\in(0,1)$, there is $\ell\geq 1$, and $q\in(0,1)$ such that for any $\Lambda\Subset\mathbb{Z}^{2}$, $t\in T$, $\Delta_{1},\Delta_{2}\subset\Lambda$, and $\xi,\xi^{\prime}\in\Omega_{\Delta_{1}}$ with $\nu_{\Lambda}^{t}(\sigma_{\Delta_{2}}=\xi)\nu_{\Lambda}^{t}(\sigma_{\Delta_{2}}=\xi^{\prime})>0$,
@@
[…]
This result is obtained from a slightly finer upper bound, but I do not believe that the finer result has a real added value compare to this one, and this statement is easier to digest. I also believe that a higher dimensional version of this result should hold (indeed: there the lack of strong mixing is manifested by the percolation of information along the boundary, which potentially has arbitrarily large percolation parameter).
Strong mixing then follows for nice enough volumes: the percolation model $P_{\Lambda;\ell,\bar{p}}$ is very subcritical in the bulk and, when the system has a “nicely one-dimensional boundary”, the inhomogeneities close to the boundary cannot destroy this decay. In the applications, the case of large squares is considered as it is easy to show exponential decay there (see Lemma A.5). But the result also holds with more complicated volumes. As far as applying the result is concerned, I believe Theorem 1.1 to be more convenient than putting some a priori restriction on the class of volume considered, and stating exponential decay for that class. Volumes under consideration might vary a great deal depending on what one is trying to do. Compared to the class of volumes considered in [2], which give strong mixing for simply connected volumes in the case of planar FK percolation, the present criterion is a priori neither weaker nor stronger: indeed, it is easy to find not simply connected regions for with it is easy to establish exponential decay under $P_{\Lambda;\ell,\bar{p}}$, and therefore apply Theorem 1.1, while it is not clear (and even wrong!) that $P_{\Lambda;\ell,\bar{p}}$ has exponential decay of connectivities uniformly over simply connected $\Lambda$. For the reader interested in less simple applications of Theorem 1.1, the arguments of [2, Section 2] permit to study exponential decay of connections probability under $P_{\Lambda;\ell,\bar{p}}$ for various family of volumes.
@@ 1.6 Relaxing the mixing hypotheses
Suppose moreover that Hypotheses (Mar1), and (Mar2) with $p_{\mathrm{bulk}}$ close enough to 11 hold (independently of the value of $C_{*}\geq 0,c_{*}>0$). Then, Hypotheses (Mix1) and (Mix2) hold.
@@ 2.2 Patch exploration
Let $(\Omega,\mathcal{F},P)$ be a probability space. Let $I$ be a finite set, and $(\mathcal{X}_{x})_{x\in I}$ a collection of finite sets equipped with the discrete $\sigma$-algebra: $\mathcal{F}_{x}=\mathcal{P}(\mathcal{X}_{x})$. Set $\mathcal{X}=\bigtimes_{x\in I}\mathcal{X}_{x}$, $\mathcal{F}=\bigotimes_{x\in I}\mathcal{F}_{x}$, and let $X$ be a $\mathcal{X}$-valued random variable. Denote $X_{J}$ the restriction (projection) of $X$ to $\bigtimes_{x\in J}\mathcal{X}_{x}$. Suppose that for every $J\subset I$, and for $P$-almost every realization of $X_{J}$, $\xi$, the conditional measure $P(\cdot\,|\,X_{J}=\xi)$ is well defined. A patch exploration sampling for $X$ is a the data of
@@ 2.3 Coarse-Graining and rescaling
Let $l\geq 0$ be the parameter linked with the “Markov-type hypotheses”. It will be more convenient to work with the blocked field $X=\psi(\sigma)$. For the rest of the section, rescale everything by $2l+1$, and identify $\mathbb{L}_{l}$ with $\mathbb{Z}^{2}$, and $\mathbb{Z}^{2}$ with $(\delta_{l}\mathbb{Z})^{2}$, where $\delta_{l}=\tfrac{1}{2l+1}$. In particular, $\sigma$ lives on $(\delta_{l}\mathbb{Z})^{2}$ and $X$ lives on $\mathbb{Z}^{2}$. for $\Lambda\subset(\delta_{l}\mathbb{Z})^{2}$, introduce
@@
Note that they are related via: $\Theta_{ij}$ is the composition of $\Theta_{ji}$ with the mirror symmetry sending $\mathrm{e}_{ij}$ to $\mathrm{e}_{ij}=\mathrm{e}_{ji}$. With these notations, and for $k\in\{\pm 1,\dots,\pm r^{L}\}$ let (see Figure 4 and Figure 5)
@@
Figure 6: Coarse-graining of a finite volume $\Lambda$: the small boxes are “blocked spins” ($\sigma$ lives on the disks, $X=\psi(\sigma)$ lives on the small boxes), the orange large boxes are $\Lambda]_{\mathrm{bulk}}$, and the blue ones are $\Lambda]_{\mathrm{bnd}}$.
@@ 2.4 Good configurations and canonical paths
Suppose that (Mar1), and (Mar2) with $p_{\mathrm{bulk}}$ close enough to 11 hold. Then, there is $L_{0}\geq 1$, such that for any $L\geq L_{0}$, and any $i\in\Gamma$,
@@
Work now under the assumption that the left-to-right and top-to-bottom crossing exist. The event $\mathrm{G}_{i}^{c}$ is then that there is $j\sim i$ and $\circ\in\{-,+\}$ such that the number of $k$’s in $\{\circ 1,\dots,\circ r^{L}\}$ with $|\tilde{\mathcal{C}}\cap\varrho(i,j,k)|\neq\varnothing$ is less than $0.75\cdot r^{L}$. This second condition implies
@@
This implies that (by a union bound over the choices for $j\sim i$ and $\circ\in\{-,+\}$)
@@
Suppose that (Mar1), and (Mar2) with $p_{\mathrm{bulk}}$ close enough to 11 hold. Then, for any $L\geq 1$, and any $e\in E_{\Gamma}$, $k\in\{\pm 1,\dots\pm r^{L}\}$,
@@
Proceed as in the proof of Claim 2: compare $\tilde{\mathcal{C}}_{E_{e}^{k}}$ with a Bernoulli percolation cluster (of parameter $p$ close to one when $p_{\mathrm{bulk}}$ is close to $1$). The absence of long crossing is then equivalent to the presence of short *-crossing of closed sites, which has exponential decay for $p$ large enough ($p$ larger than $1-e^{-2}/16$ is amply sufficient). ∎
@@ 2.6 Definition of the patch exploration process
Consider the planar labelled (multi)graph $G_{\ell}=(\Gamma,E_{\ell})$ ($\ell$ for labelled) with vertex set $\Gamma$ and $2r^{L}$ edges between neighbouring sites, labelled $\pm 1,\dots\pm r^{L}$, from the “inside” to the “outside” (see Figure 7)
@@
Now, the sequence of set valued functions $\mathcal{A}_{k},k\geq 0$ will be given by the simultaneous exploration of a percolation configuration and of a spin configuration. It is given by Algorithm 1, which requires a few preparations. Consider site-edges percolation configurations on $G_{\ell}$ with all sites of $\Lambda]_{\mathrm{ext}}$ open, all edges with at least one endpoint in $\Lambda]_{\mathrm{ext}}$ open, and all edges with both endpoints in $\Gamma\setminus[\Lambda]_{\mathrm{bulk}}$ open:
@@
  • no site in $\Delta_{2}]$ is explored, and no edge with an endpoint in $\Delta_{2}]$ is ever explored.
@@
  • $\mathcal{B}$ is the union of a collection of closed self-avoiding site-edge paths, $\gamma$, which surrounds $\Delta_{2}]$ with some edges and sites in the set $\Delta]_{2}-$surrounded by $\gamma$. $\eta_{\mathcal{B}}$ is open on $\gamma$;
@@ 2.7 Concluding the proof
[…]
@@
[…]
@@ 3 Application I: Finite range Gibbsian specifications
The reader in need of more details about Gibbsian formalism is invited to have a look at [8, Chapter 2]. Let $\Omega_{0}$ be a finite set. For $A\Subset\mathbb{Z}^{2}$, let $\phi_{A}:\Omega_{0}^{A}\to\mathbb{R}$ be a collection of uniformly bounded potential functions. The Gibbs specification associated to the collection $(\phi_{A})_{A\Subset\mathbb{Z}^{2}}$ is the collection of probability kernels $(\mu_{\Lambda}^{\xi})_{\Lambda\Subset\mathbb{Z}^{2},\xi\in\Omega_{0}^{\mathbb{Z}^{2}}}$ given by, for any event $A$ with compact support,
@@
where $Z_{\Lambda}^{\xi}$ is the normalization constant making $\mu_{\Lambda}^{\xi}$ a probability measure. For these kernels to be defined, and in order to have that the one-site marginals have uniformly bounded density with respect to the counting measure on $\Omega_{0}$, one requires the following finiteness condition:
$\sup_{x\in\mathbb{Z}^{2}}\sum_{x\in A\Subset\mathbb{Z}^{2}}\|\phi_{A}\|_{\infty}<\infty.$ (11)
Such a specification is called finite range when there is $R\geq 1$ such that $\phi_{A}=0$ for any $A$ with diameter larger than $R$.
@@
The main result of this section is a new proof of [12, Theorem 1.1].
@@
The set $(\mu_{\Lambda}^{\xi})_{\Lambda\Subset\mathbb{Z}^{2},\xi\in\Omega_{0}^{\mathbb{Z}^{2}}}$ is a particular case of the framework of Section 1.3. Take then $l>R$, the range of the potential. Then, the model satisfies (Mar1) with $\mathrm{Ma}_{u}=\Omega_{0}^{\Lambda_{3l+1}(u)}$, $\mathrm{Ma}_{u,\Lambda}^{\xi}=\Omega_{0}^{\Lambda_{3l+1}\cap\Lambda}$, (Mar2) with $p_{\mathrm{bulk}}=1$, and (Mar3) with $\mathrm{a}_{u},\mathrm{a}_{u,\Lambda}^{\xi}$ any elements of the sets $\mathrm{Ma}_{u},\mathrm{Ma}_{u,\Lambda}^{\xi}$.
Now, $(\mu_{\Lambda}^{\xi})_{\Lambda,\xi}$ are assumed to be weak mixing. So, by Theorem 1.2, they satisfy Hypotheses (Mix1), and (Mix2). Let then $\ell,q$ be the length and percolation parameter guaranteed by Theorem 1.1 for $p=10^{-10}$. Then, for $L$ large enough compared to $\ell$ (depending on $q$), it is an exercise to show that the percolation models $P_{\Lambda;\ell,\bar{p}}$ have exponential decay of connectivity uniformly over $\Lambda$ a large enough square (see remark 3.1 and Lemma A.5). ∎
@@
The statement is in fact more general: Theorem 1.1 compares mixing with connections in a (very) subcritical Bernoulli percolation with inhomogeneities. A sufficient condition for these model to have exponential decay should be that the inhomogeneity is "essentially one-dimensional" below a certain scale (e.g.: it is (well approximated by) a union of straight lines with large distances between intersection points, and lower bounded angles between lines). Obtaining optimal conditions under which such inhomogeneous site/block percolation has exponential decay seems like an interesting open problem. Handling “macroscopically one dimensional” inhomogeneities can be done as in [13, Lemma 3.2]. An argument which applies in the large squares case of Theorem 3.1 is given in Lemma A.5.